### v1

# Improving Dialog Systems with Natural Language Feedback and Preferences

Abstract
Natural language (NL) feedback contains rich information about the user experience. Existing studies focus on an instance-level approach, where feedback is used to refine specific examples, disregarding its system-wide application. This paper proposes a general framework for unlocking the system-level use of NL feedback. We show how to use feedback to formalize system-level design decisions in a human-in-the-loop-process – in order to produce better models. In particular this is done through: (i) metric design for tasks; and (ii) language model prompt design for refining model responses. We conduct two case studies of this approach for improving search query generation and dialog response generation, demonstrating the effectiveness of the use of system-level feedback. We show the combination of system-level feedback and instance-level feedback brings further gains, and that human written instance-level feedback results in more grounded refinements than GPT-3.5 written ones, underlying the importance of human feedback for building systems.
1 Introduction
When users interact with a machine learning system, they either actively or passively provide feedback to the system developers. Feedback can come in various forms, including binary ratings (Arora et al., 2022), preference feedback (Stiennon et al., 2020) and natural language (NL) feedback (Hancock et al., 2019; Scheurer et al., 2022a). Among them, NL feedback is the most general due to its free-form nature unlike the other types of feedback which have a small set of fixed options from which a user selects. It is therefore important to study how to make full use of NL feedback to enhance a system.
Existing studies on NL feedback can be summarized as belonging to one of the following two main approaches. In the first approach, feedback is used as an auxiliary target in addition to the original task, just like in multitask learning (Weston, 2016; Hancock et al., 2019; Xu et al., 2022b). In the other approach, per-instance feedback is used to modify the original output. The system can then either be fine-tuned using the new output (Tandon et al., 2022; Scheurer et al., 2022b) or undergoes an iterative process of self-critique and self-refinement at inference time (Madaan et al., 2023; Chen et al., 2023b). One common limitation of these studies is that they ignore the system-level use of NL feedback, as they only focus on using instance-level learning where each feedback only serves the specific instance for which it was received. Furthermore, they often assume the availability of feedback for each and every example, which is not practical in real world scenarios, where feedback is often sparse.
This paper asks the following research question: Can we aggregate instance-level natural language feedback to make system-level design decisions that improve language generation systems? We take the first step toward answering this question by proposing a general framework for aggregating instance-level NL feedback. A set of criteria (i.e., system-level feedback) are first derived from instance-level NL feedback through a human-in-the-loop process involving clustering and summarization. Those criteria are then used to design instruction-following language model prompts to refine (i.e., correct) all examples, and are used as references for newly designed metrics that align with users’ needs.
We conduct two case studies of the proposed framework on information-seeking dialog tasks (Glaese et al., 2022) where we improve both the query generator and the response generator of an Internet-augmented dialog system. The experimental results point to the effectiveness of system-level feedback. Our contributions can be summarized as follows.
  • We propose a new methodology that derives system-level feedback from instance-level feedback, which can guide text generation refinement.
  • We show how human experts can design quantifiable metrics for measuring different aspects of an information-seeking dialog system using system-level feedback.
  • We demonstrate that when combining system-level feedback and instance-level feedback for prompt design, the resulting refinements are more helpful for system training w.r.t. the designed metrics above.
  • We show the importance of human NL feedback by comparing it to model (GPT-3.5)-generated feedback in the response refinement process. We find that human feedback results in refinements that are more grounded, and are easier for the systems to learn from.
Figure 1: Our proposed framework for incorporating natural language feedback into model design at the system level. A human-in-the-loop process is used to derive criteria from natural language feedback, which is in turn used to design language model prompts to refine model responses, and to define metrics to measure these improvements. In the figure: $q$: query, $r$: response, $f$: feedback, $r^{\prime}$: refinement, $m(\cdot)$: metric function.
2 Related Work
2.1 Dialog Systems
With the rapid development of large pre-trained language models (Brown et al., 2020; Zhang et al., 2022), dialog systems are advancing at an incredible speed, developing all kinds of techniques, including multi-session memory (Xu et al., 2022a), search engine support (Komeili et al., 2022), etc. Recently, the emergence of ChatGPT has drawn great attention from both the NLP research community and the public at large. Nowadays, intelligent dialog agents have become an essential part of people’s productivity, such as brainstorming (Zhang et al., 2023b), essay polishing (Buruk, 2023), code writing (Haensch et al., 2023), data labeling (Gilardi et al., 2023), etc. Further, its potential in an even wider range of applications is still waiting to be unlocked.
2.2 Learning from Human Feedback
As language models begin to be used in people’s daily life, how to align them with human needs and make them more useful to users becomes an important research topic (Askell et al., 2021). As a result, researchers have been working on utilizing various kinds of human feedback, including preference feedback (Stiennon et al., 2020; Ouyang et al., 2022a), binary feedback (Li et al., 2019; Arora et al., 2022; Adolphs et al., 2022), natural language feedback (Weston, 2016; Li et al., 2017; Hancock et al., 2019; Saunders et al., 2022; Scheurer et al., 2022a), and so on. So far, the utilization of natural language feedback is relatively less explored, and the methods for using language feedback are also limited in the literature, focusing mainly on instance-level feedback where each instance receives its own feedback (Scheurer et al., 2022a; Chen et al., 2023a; Scheurer et al., 2023). In this work, we propose a general framework for deriving system-level feedback from a set of instance-level feedback, and show the effectiveness of system-level feedback alone and its complementarity with instance-level feedback.
3 Methodology
3.1 Problem Formulation
Assume we have (1) a text generator $P_{\theta}(r|q)$ that takes in a query $q$ and generates a response $r$, (2) a text refiner $P_{\phi}(r^{\prime}|r,q,c)$ that generates a refinement $r^{\prime}$ given the original response $r$, the query $q$, and some natural language described criteria $c$ that explains what makes a good output text, (3) a quality checker $Q(q,r)$ that decides whether $r$ is a satisfactory response given the query $q$. During the deployment of $P_{\theta}(r|q)$, for some unsatisfied responses $\mathcal{R}_{n}=\{r_{1},\cdots,r_{n}\}$, we collect natural language feedback for each of them $\mathcal{F}_{n}=\{f_{1},\cdots,f_{n}\}$. Our goal is to use $\mathcal{F}_{n}$ to improve $P_{\theta}(r|q)$ by updating the parameters $\theta$. In our setting, we assume that the text refiner and quality checker are given. They can either be based on general large language models like GPT-3 (Scheurer et al., 2022a) or specifically fine-tuned models (Shi et al., 2022).
3.2 Proposed Framework
Our proposed framework is shown in Figure 1. There are four steps within this framework.
Derive criteria from feedback
By deploying the text generator $P_{\theta}(r|q)$ to users, feedback $\mathcal{F}_{n}$ is collected for some responses $\mathcal{R}_{n}$. A clustering algorithm is then run (e.g., $k$-means clustering (Hartigan and Wong, 1979)) to cluster NL feedback. These clusters are used to identify common issues that can be potentially rectified. Next, a human-in-the-loop approach is used, where human experts derive a set of criteria $c$ for what constitutes a good response from those clusters. These criteria are written in natural language in a way that can be used by the text refiner as part of the input (prompt). This process is similar to prompt engineering in large-scale language models (Liu et al., 2023), where the NL feedback is used to help formalize the prompt engineering process. After deriving a set of criteria, the human experts then conduct criteria-guided metric design and design reasonable metrics $m_{1}(\cdot),\cdots,m_{k}(\cdot)$ to measure a number of aspects that users care about given $c$.
Construct refinement training data
To improve the text generator, its good behavior is reinforced, and bad behavior is discouraged. This is done through preparing training data $\mathcal{D}$ with samples served for both purposes. If a sample $(q_{i},r_{i})$ satisfies $Q(q_{i},r_{i})=1$, then it is added to $\mathcal{D}$ directly to reinforce good model behavior. Otherwise, we will use the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ to generate a refinement $r_{i}^{\prime}$. The text refiner uses the prompts based on the criteria $c$ derived from the previous step. If $Q(q_{i},r_{i}^{\prime})=1$, then the sample $(q_{i},r_{i}^{\prime})$ will be added to $\mathcal{D}$ to modify bad behavior.
Fine-tune the model
After collecting supervised data $\mathcal{D}$, we fine-tune the text generator $P_{\theta}(r|q)$ and update its parameters $\theta$. This data can be combined with existing data that was used to build the baseline deployed system (that did not use feedback).
Evaluate using designed metrics
Finally, we use our designed metrics to measure system performance and see how it aligns with the users’ requirements. If the process has been successful, the new system will have improved metrics $m_{1}(\cdot),\cdots,m_{k}(\cdot)$ compared to the baseline system, which did not incorporate feedback.
4 Experimental Setup: Dialog Systems
An appropriate place to study our framework §3.2 is in the deployment of dialogue systems, as NL feedback can be collected naturally from users as part of the interaction. For instance, an incorrect response to a user’s question might naturally prompt the user to provide feedback “that’s not correct” (Shi et al., 2022). Therefore, we consider two different case studies in the domain of dialog.
4.1 Dialog System Selection
In particular, we choose the Blenderbot 2 (BB2) dialog system (Komeili et al., 2022; Xu et al., 2022a) which contains the following two modules. (1) Query Generator (QG): given the dialog history, generate a search query for Internet search. (2) Response Generator (RG): given the dialog history and retrieved web search documents,111 We use Google search (https://www.google.com/) to retrieve the top five relevant documents given a search query. generate a final response. The reason for choosing this dialog system is that it provides two scenarios for us to verify the effectiveness of our framework: query generation and response generation.
4.2 Deployment Data
Group Feedback type Number Percentage
1
User suggests a search query for Internet search directly.
2715 52.87%
2
Gives specific editing suggestions (typically shorten the query or use more common words) for modifying the search query.
996 19.40%
3
Points out that the search query should use keywords instead of copying the original question and should be specific.
995 19.38%
4
Points out that the search query is not relevant to the problem.
429 8.35%
Table 1: Case study 1 (query generation): Four groups of system-level feedback derived from automatic clustering.
In this work, we focus on information-seeking dialog, which refers to the conversation between a user and a dialogue agent that aims to answer questions and provide relevant information (Glaese et al., 2022). We use the FITS dataset from Xu et al. (2022b) as our test bed for further analysis and experiments. FITS collects various feedback including binary feedback, NL feedback and gold corrections for human-model conversations between users and internet-augmented dialog systems such as Blenderbot 2 and SeeKeR (Shuster et al., 2022). We assume a setting where we collect binary feedback and natural language feedback only since it is less natural for conversationalists to provide gold corrections when something goes wrong.
4.3 Text Refiner
As we assume we are not given gold corrections, we investigate model-based methods to refine incorrect generations. With the rapid development of large-scale language models, powerful generalists like InstructGPT (Ouyang et al., 2022b) and ChatGPT222 https://openai.com/blog/chatgpt/ have emerged, capable of accepting only the description of a task (instruction) and accomplishing that task. Therefore, we can bypass the need for training a supervised text refiner by using them. In this work, we instantiate the text refiner using GPT-3.5333 We use the model gpt-3.5-turbo for our experiments. and we perform greedy decoding during inference.
4.4 Quality Checker
We use the collected binary feedback to train quality checkers for queries and final responses, respectively. Our classifier is based on the FLAN-T5444We use the flan-t5-large model. (Chung et al., 2022) model trained on 20% of the training data, using binary feedback following Shi et al. (2022). We select a threshold to ensure 80% precision for labels it predicts as positive on the validation set.
5 Case Study 1: Query Generation
Type Criteria
Non-copy
Specificity
Readability
Conciseness
Coverage
Satisfaction
Baseline
None
4.06
79.40
19.46
14.87
29.80
61.50
Baseline +Rephrase
(1) To better adapt to search engines, it is best not to copy the user’s original words directly. You can rephrase the user’s question, use some keywords for the search, and if the user mentions some abbreviations, restore them to their full names.
4.98
83.20
19.54
15.04
26.50
62.10
Baseline +Rephrase +Specificity
(1) To better adapt to search engines, it is best not to copy the user’s original words directly. You can rephrase the user’s question, use some keywords for the search, and if the user mentions some abbreviations, restore them to their full names. (2) Be accurate and specific enough to reflect the user’s needs.
5.00
84.20
18.77
14.50
28.80
63.30
Baseline +Rephrase +Specificity +Readability
(1) To better adapt to search engines, it is best not to copy the user’s original words directly. You can rephrase the user’s question, use some keywords for the search, and if the user mentions some abbreviations, restore them to their full names. (2) Be accurate and specific enough to reflect the user’s needs. (3) To be able to search for more results, you should use more simple and commonly used words.
5.08
80.80
19.53
15.97
29.40
62.40
Baseline +Rephrase +Specificity +Readability +Conciseness
(1) To better adapt to search engines, it is best not to copy the user’s original words directly. You can rephrase the user’s question, use some keywords for the search, and if the user mentions some abbreviations, restore them to their full names. (2) Be accurate and specific enough to reflect the user’s needs. (3) To be able to search for more results, you should use more simple and commonly used words. (4) Your search query should be concise. If the user asks multiple questions, you should focus on his/her first question.
4.81
80.00
19.70 16.63 35.30
62.70
Table 2: Case study 1 (query generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for query refinement.
5.1 Derive Criteria from Feedback
We collect all natural language feedback from the FITS training split to understand human preferences and derive criteria. We first use SimCSE sentence encoder555 We use the sup-simcse-roberta-large model. (Gao et al., 2021) to encode each feedback. Then we use $k$-means clustering to group all feedback related to query generation into five clusters and inspect the feedback in different clusters. Overall, we summarize the following four groups (see Table 1) of feedback types through merging clusters. By analyzing Table 1, we derive that a successful search query should (i) rephrase the user’s question while keeping important keywords, (ii) be relevant and specific, (iii) use some common words for better search coverage when possible, (iv) be concise. The exact criteria text which will be used to construct the prompt $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 2.
5.1.1 Criteria-guided Metric Design
After deriving criteria for improved query generation from the supplied feedback, we then design the following metrics to measure aspects of the quality of a search query that users indicated they care about in the feedback.666 When measuring the performance of a set of queries, for a metric defined as a fraction with a constant numerator, we take the average of the denominators of all queries on that metric and take its reciprocal to multiply the numerator. These metrics are designed with the guideline that a good search query will score highly in all metrics in a balanced way.
Non-copy rate
measures how much the search query rephrases the user’s utterance by examining $n$-gram matching. In particular, we define “Non-copy Rate” in Equation 1, which is based on BLEU-4 (Papineni et al., 2002). In the equation, $s$ is the search query and $u$ is the user question.
[…]
Specificity
measures whether the search query is specific enough to cover all the information needed to be able to search for the most relevant documents. We use GPT-3.5 to measure specificity (Fu et al., 2023) where we concatenate the dialog context and search query together and ask GPT-3.5 to judge whether the search query is specific using the chain-of-thought technique (Wei et al., 2022). A meta-evaluation shows 80% agreement between human annotators and GPT-3.5 for evaluating query specificity. Details can be found in the Appendix.
Readability
measures how readable is the search query by examining the word frequency rank (WFR)777 We use the list of most frequent 333,333 words in English provided by Kaggle: https://www.kaggle.com/rtatman/english-word-frequency of words in the search query. Intuitively we want the search query to contain some common words to enhance readability. We define “Readability” in Equation 2. In the equation, $w$ represents a word in $s$, $C$ is a large constant for rescaling so that the readability numbers will not be too small.
$\text{Readability}=\frac{C}{\text{AVG}_{w\in s}(\text{WFR}(w))}$ (2)
Conciseness
measures how concise the search query is by counting the number of words in the search query. We use the reciprocal of the length of the query to denote “Conciseness”. We also multiply by a constant 100 for rescaling to make a similar scale to other metrics.
Coverage
measures how specific vs. general a search query is by counting the number of Google search result pages. Since the number of pages is highly variable, we define this metric as a relative one. Given the same dialog context and the refined queries obtained using Table 2, only the query that gets the most search results has “Coverage” of 1, other queries will have “Coverage” of 0.
Satisfaction
measures whether the search query will satisfy the user. It is an overall metric, and we use our trained satisfaction classifier to measure the percentage of satisfied refinements.
5.2 Construct refinement training data
Prompt for query refinement with GPT-3.5
Given the dialog history, your task is to refine the original search query used to search the Internet so that the modified search query will search for documents that better match the user’s needs. You should follow the following requirements:
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory query.
[Original Query]
You should modify the original search query into the following:
Table 3: Case study 1: prompt for query refinement with GPT-3.5. [Criteria], [Dialog Context] and [Original Query] are placeholders to be filled. The underlined sentence is removed when [Criteria] is None.
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG) 9.74 16.09 14.28 9.61 16.09 10.15
[…]
Table 4: Performance of different query generators on FITS validation, test, and test unseen datasets, evaluated using standard F1 and PPL. The baseline BlenderBot 2 method is outperformed by our system-level trained query generator using both satisfied and refinement data (last row). Using only satisfied data is not as good (middle row).
Valid Test Test Unseen
NCR Spec. Read. Con. Cov. Sat. NCR Spec. Read. Con. Cov. Sat. NCR Spec. Read. Con. Cov. Sat.
BB2(QG) 32.8 40.5 22.4 32.3 50.6 4.8 18.8 34.9 14.0 34.3 50.9 8.8 22.7 37.7 15.4 32.9 50.3 3.2
[…]
Table 5: Performance of different query generators on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. NCR: non-copy rate, Spec.: specificity, Read.: readability, Con.: conciseness, Cov.: coverage. Sat.: satisfaction. The original BlenderBot 2 generator (top row) generates queries that have low satisfaction, despite scoring well in some other metrics. Our system-level trained query generator using both satisfied and refinement data (last row) provides much better satisfaction. Using only satisfied data is not as good (middle row).
Group Feedback type Number Percentage
1
Clarify his/her demand again.
3702 26.54%
2
Complain that the bot (i) does not answer the question (ii) or gives irrelevant information (iii) or asks the user to find out the answer on his or her own.
2260 16.20%
3
Point out specific search results that can answer the question.
2255 16.17%
4
Suggest that the bot should use the search results.
2130 15.27%
5
Point out that the bot’s answer is (i) factually incorrect, (ii) or not grounded on the search results.
1572 11.27%
6
Point out that the bot’s answer is not specific/accurate/complete/detailed.
1309 9.39 %
7
Point out that the bot is not confident in its answers and always begins its responses with “I am not sure” or “I don’t know”.
582 4.17%
8
Complain about repetition/rudeness in bot responses.
137 0.99%
Table 6: Case study 2 (response generation): Eight groups of system-level feedback derived from automatic clustering.
We first randomly sample 1,000 satisfied queries from the FITS training set together with their contexts to add to our supervised training data $\mathcal{D}$. Then, according to Figure 1-(B), for each unsatisfied Internet search query $r$, we go through the following three steps (1) refinement generation: use GPT-3.5 in combination with the criteria $c$ derived from §5.1 to get a refinement $r^{\prime}$. (2) Quality check: use a quality checker to check whether $r^{\prime}$ is satisfactory. (3) Collect data: add $(q,r^{\prime})$ to the training data $\mathcal{D}$ if $r^{\prime}$ is satisfactory, otherwise, discard it. We will describe the process of (1) in detail in the following section.
5.2.1 Refinement Generation
We instruct GPT-3.5 to generate refinements for 1,000 randomly sampled unsatisfied Internet search queries through criteria-based prompts, as shown in Table 3. To demonstrate the effectiveness of Figure 1-(A), we conduct ablation studies of using different criteria to prompt GPT-3.5 for query refinement. Considering our computational budget, for metrics calculated using GPT-3.5, we randomly sampled 500 dialog contexts and compare the queries resulting from different criteria.
The results are given in Table 2. It is clear that adding criteria in the prompt will shift the generation of GPT-3.5, and the performance differences are interpretable using the newly designed metrics. We make the following observations: (i) The rephrase criterion increases the non-copy rate. (ii) The relevance criterion increases the relevance metric. (iii) The readability criterion increases the readability and coverage metrics. (iv) When we add all the criteria, the refinements can achieve reasonably good performance in all our designed perspectives as well as in overall satisfaction. Therefore, when collecting training data, we use the four criteria augmented prompt for query refinement.
5.3 Fine-tuning the Model
In our experiments, we used the 400M parameter BlenderBot 2 query generator based on BART (Lewis et al., 2020). We consider two fine-tuning settings: (1) one with only the satisfied data; and (2) with both satisfied and refinement data. During training, we use the Adam optimizer (Kingma and Ba, 2015) with a batch size of 8 and learning rate of $7\times 10^{-6}$ for three epochs, and the best model checkpoint is chosen based on validation loss.
5.4 Evaluation using designed metrics
We compare the performance of the following query generators using both standard metrics and our designed metrics.
BB2(QG)
The original BB2 query generator.
[…]
Our system-level trained query generator using satisfied data only.
[…]
Our system-level trained query generator using both satisfied and refinement data.
5.4.1 Results on Standard Metrics
Results using standard metrics (following Shi et al. (2022)) are shown in Table 4. Comparing the performance of the original BB2 query generator and the one trained with satisfied data, it is clear that training the query generator with domain-specific data helps improve performance on F1 word overlap and perplexity (PPL) metrics by a large margin. In addition, adding refinement data for fine-tuning further enhances the performance of both metrics.
5.4.2 Results on Our Designed Metrics
We also report results on our designed metrics for different query generators, as shown in Table 5. It is clear that by training on satisfied data, the generated queries are more specific and accurate, while further adding refinement data for training will lead to even better specificity. The original BB2 query generator generates queries that are too short to be used to retrieve the most relevant documents regarding the user’s question. In other words, although it generates queries that perform well in terms of readability or coverage, it is still an inadequate query generator, as evidenced by the poor satisfaction of the queries it generates. Later, when we refer to “our trained query generator”, we mean the one trained using both satisfied data and refinement data.
6 Case Study 2: Response Generation
Type Criteria
Groundedness
Factuality
Helpfulness
Relevance
Confidence
Satisfaction
Baseline
(1) The modified response should be conversational in tone and no more than twenty words.
34.68
86.60
81.40
89.40
99.60
74.10
Baseline
+Groundedness
(1) The modified response should be conversational in tone and no more than twenty words. (2) If the user asks a question, you should use relevant search results to answer the user’s question correctly. Please do not let the user check out some resources on his or her own.
36.81
86.60
85.00
89.00
99.90
75.80
Baseline
+Groundedness +Relevance
(1) The modified response should be conversational in tone and no more than twenty words. (2) If the user asks a question, you should use relevant search results to answer the user’s question correctly. Please do not let the user check out some resources on his or her own. (3) Your modified response should be as concise and targeted as possible, and not include additional information the user has not asked for.
36.77
88.80
85.60
89.40
99.90
74.90
Baseline
+Groundedness +Relevance +Confidence
(1) The modified response should be conversational in tone and no more than twenty words. (2) If the user asks a question, you should use relevant search results to answer the user’s question correctly. Please do not let the user check out some resources on his or her own. (3) Your modified response should be as concise and targeted as possible, and not include additional information the user has not asked for. (4) Please be confident in your response, and don’t start your response with “I’m not sure” or “I don’t know”.
39.02
87.20
86.60 90.60 99.90 77.00
Table 7: Case study 2 (response generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for response refinement.
6.1 Derive criteria from feedback
Using a similar method to §5.1, we group all feedback related to response generation into ten clusters. Then, we summarize the following eight groups (see Table 6) of feedback types through splitting or merging different clusters. By analyzing Table 6, we derive that an improved response as indicated by users should (i) ground its answer on relevant search results, (ii) be concise and targeted, (iii) be confident in its answer. The exact criteria text which will be used to construct the prompts $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 7.
6.1.1 Criteria-guided Metric Design
After deriving criteria for response generation from feedback, we design the following metrics to measure the quality of a response as indicated by users.888 When using a following metric to measure the performance of a set of responses, we simply take the average of all responses’ scores on that metric.
Groundedness
measures how much the response utilizes the search results by examining $n$-gram matching. We define “Groundedness” in Equation 3, which is based on ROUGE-2 (Lin, 2004). In the equation, $r$ is the final response, and $d$ is a document from a set of relevant search documents $\mathcal{S}$.
$\text{Groundedness}=\max_{d\in\mathcal{S}}\text{ROUGE-2}(r,d)$ (3)
Factuality
measures whether the information in the response is supported by search documents. We use GPT-3.5 with chain-of-thought reasoning to measure factuality (Luo et al., 2023). We concatenate the search documents and response, and ask GPT-3.5 to judge if all information in the response is supported by the search documents. A meta-evaluation shows 88% agreement between human annotators and GPT-3.5 for evaluating response factuality. Details of the prompts and meta-evaluation can be found in the Appendix.
Helpfulness
measures whether the response directly answers the user’s question. Similar to the measurement of factuality, we use GPT-3.5 to measure helpfulness. The details are in the Appendix.
Relevance
measures whether the response provides pertinent information and stays on track without deviating from the subject. We again use GPT-3.5, with further details in the Appendix.
Confidence
measures whether the response is in a certain and confident tone. We use simple heuristics to gauge response confidence, counting the occurrences of “I’m not sure” and “I don’t know.” If either phrase appears, we consider the response as unconfident; otherwise, we assume the response to be confident.
Satisfaction
measures whether the response satisfies the user, similar to “satisfaction” in §5.1.1 .
6.2 Construct refinement training data
Prompt for response refinement with GPT-3.5
Given the dialog history and the unsatisfactory last response the bot gave, your task is to modify the response appropriately to keep the conversation fluent and consistent. You should follow the following requirements:
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory response.
[Original Response]
Below are some useful search results that you could use.
[Search Documents]
You should modify the original response into the following:
Table 8: Case study 2: prompt for response refinement with GPT-3.5. [Criteria], [Dialog Context], [Original Response] and [Search Documents] are placeholders to be filled.
Valid Test Test Unseen
F1 PPL F1 PPL F1 PPL
BB2(QG+RG) 25.78 9.40 28.30 7.41 22.99 7.75
SLT(QG)+BB2(RG) 26.69 8.24 28.66 6.66 24.88 7.03
[…]
Table 9: Case study 2 (response generation): performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using standard F1 and PPL. The original BlenderBot 2 system (top row) is outperformed by our system-level trained response generator using satisfied data (third row). However, using refinement data in this case (last row) did not help, see §6.4.1 for further anaslysis.
Valid Test Test Unseen
GRD Fact. Help. Rel. Conf. Sat. GRD Fact. Help. Rel. Conf. Sat. GRD Fact. Help. Rel. Conf. Sat.
BB2(QG+RG) 34.1 50.0 19.0 68.2 66.8 27.1 32.4 58.3 22.0 67.8 73.7 34.9 32.9 58.4 21.8 69.0 65.7 32.1
SLT(QG)+BB2(RG) 39.0 66.4 26.8 74.2 80.6 33.3 35.2 58.4 29.8 71.4 83.4 40.9 37.5 59.1 30.2 73.8 77.5 37.8
[…]
Table 10: Case study 2 (response generation): performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence. Sat.: satisfaction. The original BlenderBot 2 system (top row) is outperformed by our system-level trained response generator using satisfied and refinement data (last row). Using only satisfied data (third row) or only our query generator (second row) is not as good.
Similar to §5.2, we first randomly sample 1,000 satisfied responses together with their contexts to add to our training data $\mathcal{D}$. Then, we go through the following three steps: (1) refinement generation, (2) quality check and (3) collection of filtered data. We will describe (1) in detail in the following section.
6.2.1 Refinement Generation
We instruct GPT-3.5 to generate refinements for 1,000 randomly sampled unsatisfied responses through criteria-based prompts, as shown in Table 8. As in §5.2.1, we conduct ablation studies to demonstrate the effectiveness of derived criteria from feedback. The results are shown in Table 7. We make the following observations: (i) Adding the groundedness criterion will improve the groundedness metric. (ii) Adding the relevance criterion will increase helpfulness and relevance. (iii) GPT-3.5 generated refinements are always confident and rarely include sentences like “I’m not sure” or “I don’t know”. (iv) In terms of satisfaction, the best overall performance is achieved by the prompt with all the criteria added. Therefore, when collecting training data, we use the three criteria-augmented prompt for response refinement.
6.3 Fine-tuning the Model
In our experiments, we used the 400M parameter BlenderBot 2 main model as the baseline response generator. As before, we consider two fine-tuning settings: (1) using only satisfied data; and (2) with both satisfied and refinement data, following §5.3.
6.4 Evaluation using designed metrics
We compare the performance of the following systems:
BB2(QG+RG)
Original BB2 response model equipped with the original BB2 query generator. This is the baseline system we attempt to improve upon.
SLT(QG)+BB2(RG)
The original BB2 response generator equipped with our best system-level trained query generator from §5.
[…]
Our system-level trained response generator using satisfied data only; equipped with our best query generator from §5.
[…]
Our system-level trained response generator using both satisfied and refinement data; equipped with our best system-level trained query generator from §5.
6.4.1 Results on Standard Metrics
Prompt for response refinement with GPT-3.5 (with instance-level feedback)
Given the dialog history and the unsatisfactory last response the bot gave, your task is to modify the response appropriately to keep the conversation fluent and consistent. You should follow the following requirements:
[Criteria]
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory response.
[Original Response]
Below is the feedback for the bot’s unsatisfactory response.
[Feedback]
Below are some useful search results that you could use.
[Search Documents]
You should modify the original response into the following:
Table 11: Prompt for response refinement with GPT-3.5 (with instance-level feedback). [Criteria], [Dialog Context], [Original Response], [Feedback] and [Search Documents] are placeholders to be filled.
Valid Test Test Unseen
GRD Fact. Help. Rel. Conf. Sat. GRD Fact. Help. Rel. Conf. Sat. GRD Fact. Help. Rel. Conf. Sat.
[…]
Table 12: Case study for combining system-level and instance-level feedback: performance of different dialog systems on FITS validation, test, and test unseen datasets, evaluated using our designed metrics. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence, Sat.: satisfaction. The system trained with human feedback-guided refinement (middle row) is better than no feedback-guided refinement (first row) or GPT-3.5 feedback-guided refinement (last row).
Standard metrics are shown in Table 9. Overall, we make the following observations: (i) When using the BB2 response generator, the query generator we trained improves the quality of the final generated response compared to the BB2 query generator. (ii) Training the response generator on satisfied data leads to further improvements when using our best query generator. (iii) However, when training on both satisfied data and refinement data, there was no improvement over the case when we train on the satisfied data only.
The main reason for observation (iii) is due to the gold response collection process of the FITS dataset. Most of the time, the gold response is a human verified response that was actually generated by the original BB2 with a few rounds of guidance from the user. Therefore, using standard reference-based metrics will favor responses that come from the original BB2 distribution. In addition, low-quality references will lead to underestimates of model performance when using reference-based metrics (Zhang et al., 2023a).
6.4.2 Results on Our Designed Metrics
The results when using our criteria-based metrics are shown in Table 10. Overall, we make the following observations: (i) When using BB2 response generator, our trained query generator helps improve the quality of the final generated responses from all perspectives compared to the BB2 query generator. (ii) When equipped with our trained query generator, training the response generator on satisfied data leads to consistent improvements of the helpfulness metric compared to the BB2 response generator, which indicates the importance of domain-adapted training. (iii) Training the response generator on both satisfied data and refinement data improves the quality of final responses from all perspectives compared to training on satisfied data only, demonstrating the importance of refinement data, which provides valuable guidance on the situations where models are prone to make mistakes. (iv) In terms of satisfaction, the best overall performing system is the one that consists of our query generator and our response generator, both trained on satisfied data and refinement data.
7 Combining System-level Feedback and Instance-level Feedback
Previous studies (Scheurer et al., 2022b; Shi et al., 2022; Chen et al., 2023a) have shown the effectiveness of instance-level feedback in the refinement process. This raises the question of whether we can take advantage of both system-level and instance-level feedback to improve the performance of dialogue systems further. We use response generation as a case study to explore the answer to this question. In particular, for the 1,000 unsatisfied responses we selected in §6.2.1, we collect their corresponding human feedback as well as GPT-3.5 generated feedback (prompts can be found in the Appendix) for the refinement process. We then design a text refinement prompt that combines both system-level feedback and instance-level feedback as shown in Table 11. We then introduce the following three systems for comparison.
[…]
Our system-level trained response generator using both satisfied and refinement data; equipped with our best query generator from §5. The system does not incorporate instance-level feedback.
[…]
Our system-level trained response generator equipped with trained query generator. Here, the response generator is trained on both satisfied data and our refinement data (where we incorporated human-written instance-level feedback into the response refinement prompt).
[…]
Our system-level trained response and query generators, where the response generator is trained on both satisfied data and our refinement data. Here, we incorporate GPT-3.5, rather than human, generated instance-level feedback into the response refinement prompt.
7.1 Results of Adding Instance-level Feedback
The results when using our designed metrics are shown in Table 12. We observe that adding human-written feedback to the response refinement part brings improvements in the five criteria-based metrics most of the time, and increases the overall satisfaction consistently. However, adding GPT-3.5 generated feedback results in consistent degraded performance in groundedness, factuality and confidence. Those observations raise two questions: (1) What are the differences between GPT-3.5 generated feedback and human-written feedback? (2) How does (human/GPT-3.5) feedback impact the refinement of responses? We answer those questions in turn in the following sections.
7.2 Comparison of Human and GPT-3.5 Feedback
Groundedness Factuality Helpfulness Relevance Confidence Satisfaction
No feedback refinement 39.16 90.35 83.48 98.10 100.00 76.50
H-FB refinement 40.11 87.50 81.10 97.80 99.84 74.60
GPT-3.5-FB refinement 32.77 81.50 90.20 98.40 99.84 79.50
Table 13: Combining system-level and instance-level feedback: the quality of refinements obtained using no/human/GPT-3.5 feedback. Human feedback-guided refinement (middle row) is more grounded and factual, which end up being important factors for final model training, see §7.3.
It is interesting to investigate the differences between human-written feedback and GPT-3.5 generated feedback to understand why adding human feedback is more helpful than GPT-3.5 feedback. In particular, we characterize feedback from the following quantifiable perspectives.
Refinement Success Rate
measures whether the feedback is constructive for refining the bad response by calculating the percentage of refinements that are classified as satisfactory.
Verbosity
measures how verbose the feedback is by counting the average number of words within the feedback text.
Diversity
measures how diverse the language use is within a set of feedback. We use the percentage of unique words within the whole set of feedback to measure diversity.
Grammar
measures whether the feedback is grammatically correct. We use the percentage of feedback sentences that are grammatically correct to measure language proficiency.999 We use Gramformer for grammar error checking: https://github.com/PrithivirajDamodaran/Gramformer.
Figure 2: Characteristics of human-written and GPT-3.5 generated feedback in terms of refinement success rate, verbosity, diversity and grammar.
The characteristics of human-written and GPT-3.5 generated feedback are shown in Figure 2. Overall, we found that GPT-3.5 generates feedback text that is often more verbose with fewer grammar errors. However, the low language diversity of GPT-3.5 feedback also demonstrates that the feedback is too generic and not as targeted as human feedback. Through manual inspection, we also found that GPT-3.5 feedback tends to point out problems w.r.t. each criterion in a general tone, while human feedback is typically straightforward and to the point. Examples of GPT-3.5 feedback and human feedback are shown in the Appendix.
7.3 Feedback Impact on Refinements
While GPT-3.5 feedback generally leads to more frequent satisfaction with refinements (as indicated in Figure 2), the performance of the resulting dialog system trained with these refinements falls short w.r.t. all our designed metrics compared to the system trained using human feedback-driven refinements as shown in Table 12. This disparity merits additional investigation, emphasizing the need to delve deeper into the refinement characteristics beyond surface-level satisfaction. Therefore, we also evaluate the refinement quality using the designed fine-grained metrics from §6.1.1. The results are shown in Table 13. Refinements obtained using human feedback mainly stand out in terms of groundedness and factuality, which is in line with the feedback clusters in Table 6 where over 40% of the feedback suggests the bot focus more on the search results; that is, focusing more on the search results will make the refinements more grounded, which will lead to a more grounded final system (see results in Table 12). Since language models are known to hallucinate regardless of their size (Ji et al., 2023; Li et al., 2023), grounding their generations to the documents is important to ensure factuality. Hence, groundedness of refinements can play an essential role in the performance of trained models.
7.4 Advantages of Human Feedback
Overall, we find that feedback written by humans captures the main issues better than feedback written by GPT-3.5. For example, when a response does not answer a question, GPT-3.5 will say that the response is unhelpful because it does not contain the information the user wants. In contrast, feedback given by humans tends to tell the model the information in the search results that the user needs or gives additional hints so that the model can be guided to better generate the response. Therefore, although GPT-3.5 can generate seemingly informative feedback, it cannot replace human annotators at the moment.
8 Conclusion
In this paper, we have introduced a general framework for system-level use of natural language feedback. This is achieved by using a large set of instance-level feedback to derive a small set of criteria, that can be used for refinement prompt engineering and metric design. We demonstrate the effectiveness of system-level feedback by conducting two case studies, in generating queries and formulating dialogue responses. We further combine the system-level feedback and instance-level feedback in the refinement data construction process, and observe that the resulting trained response generator makes considerable improvements compared to using system-level feedback only. Finally, we explore the possibility of substituting instance-level human feedback with GPT-3.5 generated feedback. We found that human feedback stands out in its ability to capture the main issues, while GPT-3.5 generated feedback is lengthy and less focused. Future work could explore how to use this analysis to improve model-based feedback further to try to reach human level.
Appendix A Appendix
A.1 Evaluation with GPT-3.5
Since previous studies have demonstrated GPT-3’s reliability in the evaluation of aspects such as factuality (Luo et al., 2023), helpfulness (Fu et al., 2023), relevance (Fu et al., 2023), etc. We use GPT-3.5 to evaluate the following perspectives with the help of ChatEval (Yuan et al., 2023).
Query Specificity
Prompt for query specificity evaluation with GPT-3.5
You are evaluating a search query for a dialog using a specific set of standards. Below is the dialog context.
[Dialog Context]
Below is the search query.
[Query]
Below are the criteria.
Decide whether the search query is accurate and specific enough to enable retrieval of the most relevant documents on the Internet that are sufficient to answer the user’s question.
Does the search query meet the criterion? First, write out in a step-by-step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Then print only the single character "Y" or "N" (without quotes or punctuation) on its own line corresponding to the correct answer. At the end, repeat just the letter again by itself on a new line.
Reasoning:
Table 14: Prompt used to let GPT-3.5 evaluate query specificity. [Dialog Context] and [Query] are placeholders to be filled.
In particular, we use the prompt as shown in Table 14. Before applying it to measure the quality of query refinements. We manually labeled 50 search queries from the FITS training split, and each one’s specificity label was decided by three annotators through majority vote. We calculate the agreement between GPT-3.5 and human annotators, and the result is 80%.
Response Factuality
Prompt for response factuality evaluation with GPT-3.5
You are evaluating a response for a dialog using a specific set of standards. Below is the dialog context.
[Dialog Context]
Below are some search documents that may help continue this dialog.
[Search Documents]
Below is the response.
[Response]
Below is the criteria.
Determine if the information in the response can be found in one or more search documents.
Does the response meet the criterion? First, write out in a step-by-step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Then print only the single character "Y" or "N" (without quotes or punctuation) on its own line corresponding to the correct answer. In the end, repeat just the letter again by itself on a new line.
Reasoning:
Table 15: Prompt used to let GPT-3.5 evaluate response factuality. [Dialog Context], [Search Documents] and [Reponse] are placeholders to be filled.
The prompt we use is shown in Table 15. We conducted a meta-evaluation where we manually labeled 50 responses from the FITS training split, and three annotators decided on each one’s factuality label through majority vote. The agreement between GPT-3.5 and human annotators is 88%.
Response Helpfulness
Prompt for response helpfulness evaluation with GPT-3.5
You are evaluating a response for a dialog using a specific set of standards. Below is the dialog context.
[Dialog Context]
Below is the response.
[Response]
Below are the criteria.
Does the answer directly solve the question?
Does the response meet the criterion? First, write out in a step-by-step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Then print only the single character "Y" or "N" (without quotes or punctuation) on its own line corresponding to the correct answer. In the end, repeat just the letter again by itself on a new line.
Reasoning:
Table 16: Prompt used to let GPT-3.5 evaluate response helpfulness. [Dialog Context] and [Reponse] are placeholders to be filled.
The prompt we use is shown in Table 16. We conducted a meta-evaluation where we manually labeled 30 responses from the FITS training split and 20 responses from the Red Team dataset (Bai et al., 2022). Three annotators decided on each response’s helpfulness label through majority vote. The agreement between GPT-3.5 and human annotators is 84%.
Response Relevance
Prompt for response relevance evaluation with GPT-3.5
You are evaluating a response for a dialog using a specific set of standards. Below is the dialog context.
[Dialog Context]
Below is the response.
[Response]
Below are the criteria.
Is the response relevant to the topic at hand? It’s essential to recognize that the response does not need to be highly specific to the preceding question. As long as it remains focused on the topic at hand, it is considered relevant.
Does the response meet the criterion? First, write out in a step-by-step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Then print only the single character "Y" or "N" (without quotes or punctuation) on its own line corresponding to the correct answer. In the end, repeat just the letter again by itself on a new line.
Reasoning:
Table 17: Prompt used to let GPT-3.5 evaluate response relevance. [Dialog Context] and [Response] are placeholders to be filled.
The prompt we use is shown in Table 17. We conducted a meta-evaluation where we manually labeled 30 responses from the FITS training split and 20 responses from the Red Team dataset (Bai et al., 2022). Three annotators decided on each response’s relevance label through majority vote. The agreement between GPT-3.5 and human annotators is 84%.
A.2 Feedback Generation with GPT-3.5
Prompt for feedback generation with GPT-3.5
Given the dialog history, the unsatisfactory last response the bot gave, and the requirements for a good response, your task is to write detailed and constructive feedback to improve the unsatisfactory response. The requirements for a good response include the following:
(1) The response should be conversational in tone and no more than twenty words.
(2) If the user asks a question, the response should use relevant search results to answer the user’s question correctly. It should not leave the user to check out some resources on his or her own.
(3) The response should be as concise and targeted as possible, and not include additional information the user has not asked for.
(4) The response should be in a confident tone and not start with “I’m not sure” or “I don’t know”.
Below is the dialog context.
[Dialog Context]
Below is the bot’s unsatisfactory response.
[Original Response]
The feedback for the unsatisfactory response is the following:
Table 18: Prompt for feedback generation with GPT-3.5. [Dialog Context] and [Original Response] are placeholders to be filled.
Previous studies have demonstrated the capability of large language models to generate informative and useful feedback (Madaan et al., 2023; Chen et al., 2023b). Therefore, we also investigate using GPT-3.5 to generate instance-level feedback for each unsatisfied response. The prompt we use for feedback generation is in Table 18.
A.3 Examples of GPT-3.5 and Human Instance-level Feedback
We list examples of instance-level feedback written by humans and GPT-3.5 in Table 19.
Human written instance-level feedback GPT-3.5 written instance-level feedback
Good. You just need to include the reason and his age at the time of his death in addition to this for a complete answer.
The response is not relevant to the user’s question and includes additional information that the user did not ask for. A good response should focus on answering the user’s question about how Fitzgerald died, using relevant search results. It should be concise, confident, and conversational in tone.
You could tell me to take my cat to the vet when they get sick.
The response is not helpful. Please provide information on treatment options for cat rhinovirus. Use relevant search results to answer the user’s question correctly and be concise in your response. Avoid providing additional information the user has not asked for.
The bot did not list Azerbaijan or Estonia, which were included in the search results. The bot should include these two other Soviet Republics in the response.
The response is informative but could be improved. Instead of listing the republics in a long sentence, try breaking them down into bullet points for easier reading. Also, it would be helpful to mention that there were 15 republics in total, as the user requested.
Again, the bot’s "answer" doesn’t really make sense. It claims that a character, not an actor, played the part of Celie.
The response is confusing and irrelevant. Please provide a clear and concise answer to the user’s question using relevant search results. Avoid adding unnecessary information or asking unrelated questions.
The answer is right there in the results, and Charles de Gaulle isn’t correct.
The response is good, but it could be improved by being more conversational and concise. Instead of just stating the answer, try something like "Charles de Gaulle was the first French president in 1848. Is there anything else I can help you with?" This makes the response more engaging and shows confidence in the answer.
Interesting. We need to know which premier specifically, with data to back up the claim. Can we be more specific?
The response is not helpful as it does not answer the user’s question. Please provide a specific answer with relevant search results to support it. Avoid guessing or providing irrelevant information. Be confident and concise in your response.
Search about living creatures specifically.
The response is not relevant to the user’s question. Please provide a direct answer to the user’s question using relevant search results. Avoid providing additional information that the user has not asked for. Be confident in your response and avoid starting with "I’m not sure" or "I don’t know".
Actually it looks like that happened way later. Let’s go back to the Revolutionary War and how that affected Britain.
The response is not helpful as it does not answer the user’s question. A better response would be to use relevant search results to explain the reasons for the decline of the British empire, such as economic challenges, decolonization movements, and global power shifts. The response should be confident and concise, and not include additional information the user has not asked for.
Use the search results provided to formulate some answer. Mention how there are wave patterns that one meteorologist uses and got 2/3 correct, mention a tool that they use like the doppler radar which looks for things like airborne tornado debris.
The response is too general and does not answer the user’s question. Instead, use relevant search results to provide specific weather conditions that meteorologists look for to predict tornado formation. Keep the response concise and confident in tone.
One of the search results specifies that Paul Tillich was Rollo May’s mentor and friend, confirming that they did know each other.
The response should have used relevant search results to answer the user’s question. It should have been more targeted and concise, and not included additional information the user did not ask for. A better response would have been: "I’m not sure if they knew each other, but according to my search results, Tillich was a major influence on Rollo May’s work."
Why didn’t you mention CrossFit or Tabata?
The response should have used relevant search results to answer the user’s question correctly. Instead of saying "I’m not sure," the bot should have confidently provided the correct answer. Additionally, the response should have been more concise and targeted, only providing the specific exercise type that burns the most calories.
Table 19: Examples of human written and GPT-3.5 written instance-level feedback. For each row, both human feedback and GPT-3.5 feedback are written to the same unsatisfied response.